OpenStudy (anonymous):

prove for every 'n' {3^3n-2+3^6n-4+1} divide by 13

OpenStudy (anonymous):

lord this is going to be a raft of algebra because you probably have to do it by induction right?

OpenStudy (anonymous):

yea..

OpenStudy (anonymous):

i can't possibly write it all out here, and the algebra is going to suck, but the idea is this first show it is true if \(n=1\) which it is because you get \(13\)

OpenStudy (anonymous):

then replace \(n\) by \(k\) that is what you get to assume is true \[\large 3^{3k-2}+3^{6k-4}-1\] is divisible by 13

OpenStudy (anonymous):

yea and after that i did for k+1 and i stuck

OpenStudy (anonymous):

next replace \(n\) by \(k+1\) and get \[\large 3^{3k+1}+3^{6k+2}-1\]

OpenStudy (anonymous):

true.. but i hve problem to make it like the basis of the induction

OpenStudy (anonymous):

then see if you can rewrite this as \(3^{3k-1}+3^{6k-2}-1+something\)

OpenStudy (anonymous):

there i need your help man

OpenStudy (anonymous):

typo there, but you get the idea should be \[\large 3^{3k-1}+3^{6k-4}-1+something\] the something will obviously be divisible by 13, and the first part is by induction

OpenStudy (anonymous):

ok let me try with pencil and paper, i can't do it on the fly

OpenStudy (anonymous):

great! take your time.. its a lot of games between the numbers..

OpenStudy (anonymous):

almost there

OpenStudy (anonymous):

got it you still there?

OpenStudy (anonymous):

yeaa man..

OpenStudy (anonymous):

ok we are at this step \[\large 3^{3k+1}+3^{6k+2}-1\] right

OpenStudy (anonymous):

why -1?

OpenStudy (anonymous):

oh it should be \[\large 3^{3k+1}+3^{6k+2}+1\]

OpenStudy (anonymous):

it is a \(+1\) out at the end yes?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok now we want the previous exponents \(3k-2\) and \(6k-4\) so step one is to write each as \[\large 3^3\times 3^{4k-2}+3^6\times 3^{6k-4}+1\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok now factor out the \(3^3\) and insert a \(1\) in the parentheses to make it look like the previous case \[\large 3^3\times 3^{4k-2}+3^6\times 3^{6k-4}+1\] \[=3^3\left(3^{4k-2}+3^{6k-4}+1\right)-3^3+(3^6-3^3)\times 3^{6k-4}+1\]

OpenStudy (anonymous):

the \(3^6-3^3\) is probably the confusing part, or maybe not but when you factor out the \(3^3\) you have inside the parentheses \(3^3\times 3^{6k-4}\) but you need \(3^6\times 3^{6k-4}\) so you have to add back \((3^6-3^3)\times 3^{6k-4}\) at the end

OpenStudy (anonymous):

once we have \[\large =3^3\left(3^{4k-2}+3^{6k-4}+1\right)-3^3+(3^6-3^3)\times 3^{6k-4}+1\]we are pretty much done the first part is divisible by 13 by induction, as it is \(27\) times something we know is divisible by 13 it turns out luckily that \(3^6-3^3=702\) is also divisible by 13

OpenStudy (anonymous):

and also \(1-3^3=-26\) is also divisible by 13

OpenStudy (anonymous):

guess it would be better to write \[\large =3^3\left(3^{4k-2}+3^{6k-4}+1\right)-3^3+(3^6-3^3)\times 3^{6k-4}+1\] \[=\large =3^3\left(3^{4k-2}+3^{6k-4}+1\right)+702\times 3^{6k-4}-26\]

OpenStudy (anonymous):

or even \[\large =3^3\left(3^{4k-2}+3^{6k-4}+1\right)+13\times(54\times 3^{6k-4}-2)\]

OpenStudy (anonymous):

told you the algebra was going to suck, and i didn't lie

OpenStudy (anonymous):

lol yea! thanks a lot for your time . you really helped me! thx again!

OpenStudy (anonymous):

yw